1-Consider the following reaction at equilibrium. The Kc value for the reaction at 295 K is 1.20 x 10-4.

NH4HS (s) ⇌ NH3 (g) + H2S (g)

a. Calculate the Kp value at 295 K.

b. Determine the partial pressure of each gas at equilibrium.

c. Calculate the Kc value of the following reaction.

3NH3 (g) + 3H2S (g) ⇌ 3NH4HS (s)
Answer #1

1) Kp = Kc(RT)∆n

= 1.20×10-4(0.0821×295)2

= 0.0704

2) P​​​​​​2 = 0.0704

P = 0.265 atm

Partial pressure of each gas at equilibrium = 0.265 atm

3) Kc = (1/1.2×10-4)3

Kc = 5.79×1011